Sunlight Sailboats has done its pro forma balance sheet and cash budget and has determined that they will need an additional $350,000 of cash for one month. They have the following three options. Find the cost of each option.   

a. Establish a one year line of credit for $350,000 with a bank. The commitment fee will be .5% of the unused portion of the line of credit. The interest charge on the used portion will be 12 percent per year. Assume that the funds are only needed and will only be used for one month.

b. Forego the July trade discount of 3/10, net 40 on $350,000 of accounts payable.

c. Take out a discount loan for the month the financing is needed. The loan will have an interest rate of 14%.

d. Which option do you recommend?
Answer #1
a.
Total Interest - Line of Credit
Interest on used portion for 1 month $   3,500.00
=350000*12%*1/12
Interest in unused portion for 11 months $   1,604.17
=350000*.5%*11/12
Total Interest $   5,104.17
b.
Loss due to Cash discount not availed
Cash Discount not availed on accounts payable $ 10,500.00
=350000*3%
Total Interest $ 10,500.00
c.
Total Interest - Discount Loan
Interest on 1 month $   4,083.33
=350000*14%*1/12
Total Interest $   4,083.33
Hence the best option is c. to take a discount loan at 14%.